# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v2

@@ 8.2.1. Hardware Layer
The deployment of billions of cameras in roads, factories, and public places not only provides diverse application scenarios but also offers ample data support for training large-scale NSVAD models. Next-generation communication devices and image processing units can provide communication and computational support for building large-scale NSVAD systems, making it feasible for cloud-based global model learning in end-cloud collaborative architecture. Additionally, emerging sensing devices such as thermal imaging, motion cameras, and event cameras will expand the application potential of NSVAD in scenarios like military and sports. Therefore, we believe that the advancement and innovation of hardware devices strongly drive the development and application of NSVAD technology, and developing deployable NSVAD systems for new types of sensors and large-scale IoT systems will become a trend.
@@ 8.2.2. System Layer
The system layer aims to bridge the hardware and algorithm layers, providing interfaces for model deployment on terminal devices and supporting the collaborative optimization of computation and communication execution in NSVAD systems. Previous NSVAD research focused on algorithm design while neglecting system layer development for a long time. The end-cloud collaborative framework introduced in Section 7.3 explicitly indicates the importance of the system layer in promoting the video anomaly detection technology in large-scale IoT and mobile internet applications and designs the first NSVAD system layer for end-cloud collaborative detection and resource allocation. With the further development of edge artificial intelligence and mobile communication networks, new solutions will be sought for communication strategy optimization, computation offloading, distributed model aggregation, and interface flexibility faced by the system layer, promoting NSVAD towards integration and intelligence.
@@ 8.2.3. Algorithm Layer
Combining the latest advancements and domain concerns, we believe the development opportunities for algorithms include: 1) the introduction of large-scale heterogeneous datasets; 2) the adaptability transfer of efficient representation learning methods; 3) innovative combinations of emerging artificial intelligence tasks with NSVAD; and 4) assistance from implicit knowledge in large models (Han et al., 2024a). Specifically, expanding business scenarios from smart transportation and modern factories provide rich data sources for VSVAD, making it possible to train large-scale models. The rise of online video networks and streaming platforms provides additional data modalities beyond images, such as audio, subtitles, and language, aiding in exploring complex anomaly clues. Furthermore, the development of virtual data engines allows easy emulation of rare anomaly events and provides pixel-level annotations. The rarity and diversity of anomalies and the difficulty of collecting cross-scene videos will no longer be bottlenecks for NSVAD model development. Rich data modalities and volume will drive the development of multimodal and cross-scene NSVAD models and are expected to foster supervised interpretable research routes. The development of deep learning will free NSVAD algorithms from manual features and enable them to extract spatial-temporal features end-to-end and model video normality, while the continuous advancement of deep neural networks (such as attention, Transformer, masked autoencoders) and emerging representation learning methods (such as video self-supervised learning and causal representation learning) will further improve NSVAD’s learning capability. The advancement of emerging artificial intelligence technologies will provide references for addressing specific concerns of NSVAD. For example, unsupervised NSVAD recovers from label-agnostic data biases in complex environments, often encountering significant performance degradation when dealing with diverse common events, a problem long studied in domain generalization tasks. We believe the combination of domain generalization and NSVAD research will help address the negative impact of data bias in unsupervised solutions. Data security and privacy protection have always been concerns for users and researchers when building deployable NSVAD systems for large-scale IoT. Federated learning will provide feasible solutions. The rise of large models will propel generative artificial intelligence research to a climax and give birth to numerous phenomenon-level applications, reshaping industries, and NSVAD is no exception. We believe that large models contain implicit knowledge related to anomalous events, which is crucial for understanding the fundamental differences between normal and abnormal and developing interpretable NSVAD models.
@@ Acknowledgements.
This work is supported in part by the China Mobile Research Fund of MOE under Grant KEH2310029, NSFC under Grant 62250410368, and China Scholarship Council (File No. 202306100221). The authors would like to express their gratitude to the anonymous reviewers for their insightful comments and valuable suggestions.

### v3

@@ 7.2.1. Hardware Layer
The deployment of billions of cameras in roads, factories, and public places not only provides diverse application scenarios but also offers ample data support for training large-scale NSVAD models. Next-generation communication devices and image processing units can provide communication and computational support for building large-scale NSVAD systems, making it feasible for cloud-based global model learning in end-cloud collaborative architecture (Ju et al., 2024; Teng et al., 2024). Additionally, emerging sensing devices such as thermal imaging, motion cameras, and event cameras will expand the application potential of NSVAD in scenarios like military and sports. Therefore, we believe that the advancement and innovation of hardware devices strongly drive the development and application of NSVAD technology, and developing deployable NSVAD systems for new types of sensors and large-scale IoT systems will become a trend.
@@ 7.2.2. System Layer
The system layer aims to bridge the hardware and algorithm layers, providing interfaces for model deployment on terminal devices and supporting the collaborative optimization of computation and communication execution in NSVAD systems. Previous NSVAD research focused on algorithm design while neglecting system layer development for a long time. With the further development of edge artificial intelligence and mobile communication networks, new solutions will be sought for communication strategy optimization, computation offloading, distributed model aggregation, and interface flexibility faced by the system layer, promoting NSVAD towards integration and intelligence.
@@ 7.2.3. Algorithm Layer
Combining the latest advancements and domain concerns, we believe the development opportunities for algorithms include: 1) the introduction of large-scale heterogeneous datasets; 2) the adaptability transfer of efficient representation learning and reinforcement learning methods (Liu et al., 2024b, 2023e, 2022g); 3) innovative combinations of emerging artificial intelligence tasks with NSVAD; and 4) assistance from implicit knowledge in large models (Han et al., 2024a). Specifically, expanding business scenarios from smart transportation and modern factories provide rich data sources for VSVAD, making it possible to train large-scale models. The rise of online video networks and streaming platforms provides additional data modalities beyond images, such as audio, subtitles, and language, aiding in exploring complex anomaly clues. Furthermore, the development of virtual data engines allows easy emulation of rare anomaly events and provides pixel-level annotations. The rarity and diversity of anomalies and the difficulty of collecting cross-scene videos will no longer be bottlenecks for NSVAD model development. Rich data modalities and volume will drive the development of multimodal and cross-scene NSVAD models and are expected to foster supervised interpretable research routes. The development of deep learning will free NSVAD algorithms from manual features and enable them to extract spatial-temporal features end-to-end and model video normality (Zhou and Wang, 2024; Liu et al., 2024g; Zhou et al., 2024b; Liu et al., 2024f), while the continuous advancement of deep neural networks (such as attention, Transformer, masked autoencoders) (Zhou and Wang, 2022; Liu et al., 2022i; Zhou et al., 2019a; Liu et al., 2022h) and emerging representation learning methods (such as video self-supervised learning and causal representation learning) will further improve NSVAD’s learning capability. The advancement of emerging artificial intelligence technologies will provide references for addressing specific concerns of NSVAD. For example, unsupervised NSVAD recovers from label-agnostic data biases in complex environments, often encountering significant performance degradation when dealing with diverse common events, a problem long studied in domain generalization tasks. We believe the combination of domain generalization and NSVAD research will help address the negative impact of data bias in unsupervised solutions. Data security and privacy protection have always been concerns for users and researchers when building deployable NSVAD systems for large-scale IoT. Federated learning will provide feasible solutions. The rise of large models will propel generative artificial intelligence research to a climax and give birth to numerous phenomenon-level applications, reshaping industries, and NSVAD is no exception. We believe that large models contain implicit knowledge related to anomalous events, which is crucial for understanding the fundamental differences between normal and abnormal and developing interpretable NSVAD models.
@@ Acknowledgements.
This work was supported by the National Key Research and Development Program of China under Project No. 2024YFE0200700 and Subject No. 2024YFE0200703. This work was also supported in part by the Specific Research Fund of the Innovation Platform for Academicians of Hainan Province under Grant YSPTZX202314, in part by the Shanghai Key Research Laboratory of NSAI and the Joint Laboratory on Networked AI Edge Computing, Fudan University-Changan. This work was also supported in part by the China Mobile Research Fund of MOE under Grant KEH2310029 and in part by the NSFC under Grant 62250410368. The authors sincerely thank Liangyu Teng, Yuntian Shi, and Hao Yang from the Fudan University for their help in revising this article. The authors would like to express their gratitude to the anonymous reviewers for their insightful comments and valuable suggestions.